# How transparent is graphene? A surface science perspective on remote epitaxy

### v1

@@ Abstract
Remote epitaxy is the synthesis of a single crystalline film on a graphene-covered substrate, where the film adopts epitaxial registry to the substrate as if the graphene is transparent. Despite many exciting applications for flexible electronics, strain engineering, and heterogeneous integration, an understanding of the fundamental synthesis mechanisms remains elusive. Here we offer a perspective on the synthesis mechanisms, focusing on the foundational assumption of graphene transparency. We highlight challenges for quantifying the strength of the remote substrate potential that permeates through graphene, and propose Fourier and beating analysis as a bias-free method for decomposing the lattice potential contributions from the substrate, from graphene, and from surface reconstructions, each at different frequencies. We highlight the importance of graphene-induced reconstructions on epitaxial templating, drawing comparison to moiré epitaxy. Finally, we highlight the role of the remote potential in tuning surface diffusion and adatom kinetics on graphene. Tuning the surface diffusion length is crucial in navigating the competition between remote epitaxy and defect-seeded mechanisms like pinhole epitaxy.
@@ I Introduction
Remote epitaxy [1] is a transformative approach for synthesizing single crystalline membranes and lattice mismatched heterostructures, for application in flexible electronics [2], discovery of new properties via strain engineering [3, 4, 5, 6, 7], lattice mismatched heteroepitaxy with reduced dislocation densities [8, 9, 10], and integration of dissimilar materials [11]. In this approach, thin films are grown on a graphene-covered single crystalline substrate. Under the right conditions, a single crystalline film can be produced which has epitaxial registry to the substrate rather than to graphene. Several reviews have focused on applications [12, 13, 14], the processing steps required for making, transferring, and performing epitaxy on top of graphene-covered surfaces [15, 14, 16], and comparison with different types of epitaxial growth modes [17, 18].
@@
Here we highlight some of the outstanding questions regarding graphene transparency, and their impacts on thermodynamics and kinetics of remote epitaxy:
@@ II How strong is the remote lattice potential?
To date, the field has relied on several proxies for the remote lattice potential. In the pioneering work on remote epitaxy, Kim et. al. [1] computed the charge density as a function of distance for GaAs slabs that are separated by a vacuum gap, using density functional theory (DFT). They found that for slabs separated by up to 9 Å, there is a finite charge density in the middle of the vacuum gap. This calculation provided the first theoretical suggestion that the lattice potential of the substrate may extend significantly beyond the Å typical equilibrium spacing between atoms. It should be noted, however, that within the DFT surface science community, the emergence of charge within the vacuum region is considered a sign that the vacuum region is not large enough to create a truly “isolated” surface. While the calculations in question are trying to demonstrate that interaction between a substrate and film can take place across larger vacuum distances, this ignores the periodic nature of the structure. I.e., the DFT calculation represents an infinite GaAs material with repeating vacuum regions in the middle. It is not clear from these calculations that this charge density interaction through the vacuum would persist if the film and substrate were truly isolated. Furthermore, the periodic nature of these calculations implies both the film and substrate are of the same size, which at best describes a minority of homoepitaxial cases. The finite slab size would need to be tested for convergence to ensure that both ends of the slab could behave independently to truly mimic a substrate and film with the ability to differ.
@@
A second challenge is that the raw potential maps contain more than one periodicity: the substrate lattice, the graphene lattice, and a longer-range supercell periodicity from relaxation, such that
To distinguish of the remote substrate, Ref. [24] used the following subtraction and cropping procedure: First, they subtract the short period potential for isolated graphene (), resulting in the maps of Fig. 1(c,d). Next they crop the longer range fluctuation of the relaxed supercell (): for the profile above graphene/GaN (0001) (Fig. 1(d,f)), they focus on the line cut from atom 3 to atom 1, crop away the rest of the supercell, then subtract a linear background. The energy difference along the path from atom 3 to atom 1 is then used as for the substrate, and this selected profile is periodically repeated in the line cut shown in Fig. 1(h).
@@
A recent analytical model by Kawasaki and Campbell further suggests that once free carrier screening from graphene is included, the remote substrate potential and the graphene potential have similar magnitude [34] (Fig. 3), and thus interference with graphene cannot be ignored. This model starts from a Morse pair interaction potential (rather than electrostatics), which includes covalent and van der Waals interactions between film and substrate, plus a van der Waals spacer layer modeled by a Lennard-Jones potential. The inclusion of free carrier screening for typical graphene carrier densities significantly attenuates the remote potential, such that for a typical graphene densities of order cm-2, the screened potential is of order 10 meV, similar to graphene. The advantages of this approach are that the Morse potential describes a pairwise bonding interaction, rather than the purely electrostatic interaction computed by DFT, and is defined by simple, interpretable parameters that are well benchmarked by experiments and theory.
@@ III To what extent does graphene interact with a substrate or with a film?
Buffer graphene on Si-face SiC (0001), which forms a reconstruction, is a canonical example. A scanning tunneling microscopy (STM) image of this surface is shown in Fig. 4(e). Here the buffer carbon layer is not fully graphitic. Instead, core level spectroscopy reveals that it has mixed hybridization with and character [30], strong bonding to the SiC substrate, out-of-plane dangling bonds, Å buckling in the graphene, and both in-plane and out-of-plane distortions in the first layer of SiC [30, 41, 42, 43]. This deviation from graphitic character also opens a bandgap in buffer graphene [31, 44]. The bandgap is important because it reduces the free carrier screening of the substrate potential, compared to the screening from metallic graphene. Thus, the remote potential through a semiconducting buffer graphene is expected to be larger than the remote substrate potential through metallic graphene. However, the buckled buffer layer, with its dangling bonds, may introduce larger local potential fluctuations than flat graphene, potentially dominating over any remote field effects from the substrate.
@@
The large structural distortion of buffer graphene has a significant impact on adatom sticking and film growth. Figure 5(b) shows an STM image of Co atoms deposited on a surface that contains regions of buffer graphene and regions with a second “epitaxial” layer of graphene on top of the buffer [45]. Co clusters stick with higher probability to buffer graphene, than to epitaxial graphene. This preferred sticking on buffer graphene likely arises from the dangling bonds in hybridized buffer graphene, compared to the more idealized character of epitaxial graphene. Presumably this change in bonding character would be relevant for the growth of remote epitaxial films on SiC buffer layer, indicating a large role for reconstruction in these systems.
@@
Graphene on Ge (110) also reconstructs, into an apparent “” pattern (Fig. 4(f)). This reconstruction is characterized by lateral and out of plane reconstructions of the top two Ge layers, with minimal distortions in the graphene itself [37, 46, 47]. Here the graphene retains its semimetallic character. Note while this reconstruction is commonly referred to as “”, its true periodicity is defined by , where are the vectors of the surface reconstruction and are the Ge(110) surface unit cell vectors [37].
@@
Finally, even graphene on metallic surfaces induces reconstructions. Fig. 4(g) shows a STM image of graphene on Ir (111). The 10% lattice mismatch between graphene and Ir produces a supercell with nm periodicity, corresponding to Ir unit cells and graphene unit cells. Photoemission and scanning probe measurements reveal that the graphene is buckled by Å, due to the periodic modulation in Ir-C atomic alignment, which changes the Ir - C hybridization strength [48, 49, 38]. Moiré-driven reconstructions have also been observed on gr/Ru(0001)[50, 51, 52], gr/Rh(111)[53, 54], gr/Ni(111)[55, 56], gr/Pt(111)[57], and gr/Cu(111)[58, 59, 60] with interaction strengths spanning from strong chemisorption (e.g., Ni, Rh) to weak van der Waals bonding (e.g., Cu, Pt) [49, 48].
@@
Epitaxial ordering on graphene/metal moiré templates lies in apparent contrast to the polar bonding hypothesis for remote epitaxy. Recall that the electrostatic potential analysis by Kong et. al. suggested the remote potential fluctuations should be strongest for materials with polar bonding [24]. In contrast, Ir (111) has metallic bonding. For Ir on graphene/Ir (111) the ordering is thought to be driven by the reactivity of Ir to the HCP sites within the moiré cell [38], rather than remote epitaxy through transparent graphene. Further studies are required to connect the ordering of adatom clusters on graphene/substrate moiré templates, which have more periodicity of a few nanometers, to the continuous films of remote epitaxy that have lattice constant of a few Angstrom.
Furthermore, these reconstructions complicate the initial mechanistic picture of remote epitaxy that the graphene layer allows the film periodicity to less strongly match the substrate periodicity. Which lattice would the film be interacting with: the bulk substrate or the graphene induced reconstruction (or both)? If the insertion of graphene removes the film’s lattice matching requirement with the bulk substrate, but replaces this with a lattice matching requirement with the resulting reconstruction, then remote epitaxy is arguably just an extension of regular epitaxial principles to a new surface induced by the addition of graphene.
@@ IV How do graphene/substrate interactions tune adatom kinetics?
Just as the underlying substrate is expected to tune the remote lattice potential, the substrate is also expected to tune adatom diffusion on graphene-covered surfaces. Growth kinetics have been underexplored in the context of remote epitaxy, but have been considered more extensively in the surface science literature. Fig. 6 summarizes the DFT computed surface diffusion barrier heights for several metal adatoms on graphene-covered metal surfaces [64, 65, 66, 67, 68, 69, 70, 71]. In Fig. 6(a) we plot versus the measured or computed topographical corrugation height of graphene on these surfaces, since this corrugation is a proxy for the graphene-substrate interaction strength [49]. As the substrate is changed, for Au adatoms systematically increases going from bare graphene ( Å)[72] to graphene/Cu (111) ( Å),[73] to buffer graphene/SiC ( Å)[72] to graphene/Ru(0001) ( Å).[74] For a fixed graphene/Ru(0001) substrate and varying the adatom (Fig. 6(b)), the barrier height also changes significantly and follows a general trend of valence electron count of the adatom. These data suggest that kinetics, which are tunable via the graphene/substrate interface, may play a significant role in tuning the growth mechanisms and structure of films grown on graphene-covered surfaces, in addition the the thermodynamic conditions discussed in the previous sections. “Remote” tuned diffusion barriers go beyond the typical kinetic tuning knobs of growth rate and sample temperature.
@@
For epitaxy on graphene-covered surfaces, the kinetic regimes are not well known, although one limiting case has been studied. For transferred graphene-covered surfaces, where pinholes are the primary defect, Manzo et. al. showed that when the diffusion length is greater than the spacing between pinholes , nucleation occurs preferentially at pinholes where there is direct bonding to the substrate [19]. Note that where is the diffusion time. This selective nucleation is followed by lateral coalescence of a continuous film. Since the pinholes are typically only a few nanometers in diameter, films grown by the pinhole mechanism can be exfoliated, and thus share many of the macroscopic hallmarks that are expected for films that grow by a remote epitaxy mechanism. Further studies are required to investigate how tuning surface diffusion, and tuning both the identity and concentration of different types of defects, tunes the growth mode on graphene-covered surfaces. While a pinhole mechanism dominates when there is a high concentration of pinholes or larger openings in graphene [77, 78, 79, 80, 19, 81, 82], perhaps other sites such as graphene-covered atomic step edges or local variations in the graphene/substrate registry could be the nucleation sites for a true remote epitaxy mechanism (Fig. 6(c)).
V Outlook

### v2

@@ Abstract
Remote epitaxy is the synthesis of a single crystalline film on a graphene-covered substrate, where the film adopts epitaxial registry to the substrate as if the graphene is transparent. Despite many exciting applications for flexible electronics, strain engineering, and heterogeneous integration, an understanding of the fundamental synthesis mechanisms remains elusive. Here we offer a perspective on the synthesis mechanisms, focusing on the foundational assumption of graphene transparency. We identify challenges for quantifying the strength of the remote substrate potential that permeates through graphene, and propose Fourier and beating analysis as a bias-free method for decomposing the lattice potential contributions from the substrate, from graphene, and from surface reconstructions, each at different frequencies. We highlight the importance of graphene-induced reconstructions on epitaxial templating, drawing comparison to moiré epitaxy. We highlight the role of the remote potential in tuning surface diffusion and adatom kinetics on graphene, which are crucial for navigating the competition between remote epitaxy and defect-seeded mechanisms like pinhole epitaxy. In light of this weak remote potential, we re-evaluate the current state-of-the-art experimental evidence, highlighting why it remains challenging to experimentally validate a “remote” epitaxy mechanism that cannot be explained by alternatives, such as pinhole-seeded epitaxy or serial van der Waals epitaxy. We end with one experimental example that, to out knowledge, cannot be explained by competing mechanisms: a different long-range epitaxial relationship for GdPtSb films grown on graphene/sapphire, compared to direct epitaxy on sapphire. We suggest for future experiments that directly measure the remote potential and impact of tuneable growth kinetics.
@@ I Introduction
Remote epitaxy[1] is an approach for synthesizing single crystalline films that are semi-decoupled from their substrates, to circumvent the lattice [2, 3] and chemical [4] mismatch challenges of traditional epitaxy. In this approach, thin films are grown on a graphene-covered single crystalline substrate. Under the right conditions, a single crystalline film can be produced which has epitaxial registry to the substrate rather than to graphene. To date, this concept has been applied to the synthesis of III–V and III–N semiconductors [1, 5, 6, 7, 8], transition metal oxides[9, 10], halides[11, 7], and intermetallic compounds[12, 13, 14, 15]. Applications of these heterostructures include membrane exfoliation for flexible electronics [16, 17], discovery of new properties via strain engineering in membranes [14, 12, 18, 19, 20], lattice mismatched heteroepitaxy with reduced dislocation densities [6, 7, 8], integration of dissimilar materials [11, 18], e.g., for stacking multi color (bandgap) micro-LED arrays [21], blocking diffusion across film/substrate interfaces[22], and reuse of expensive substrates[23]. Several reviews have focused on applications [24, 13, 25], the processing steps required for making, transferring, and performing epitaxy on top of graphene-covered surfaces [26, 25, 27], and comparison with different types of epitaxial growth modes [28, 29].
@@
Here we highlight some of the outstanding questions regarding graphene transparency, their impacts on thermodynamics and kinetics of remote epitaxy, and why it remains an outstanding challenge to definitively demonstrate a remote epitaxy mechanism that cannot be explained by alternatives like pinholes [30] or serial van der Waals epitaxy:
@@
Section V: Given the weak remote substrate potential, how can we explain existing experiments on “remote” epitaxy?
Section VI: Outlook and call for new experiments and theory.
@@ II How strong is the remote lattice potential?
To date, the field has relied on several proxies for the remote lattice potential. In the pioneering work on remote epitaxy, Kim et. al. [1] computed the charge density as a function of distance for GaAs slabs that are separated by a vacuum gap, using density functional theory (DFT). They found that for slabs separated by up to 9 Å, there is a finite charge density in the middle of the vacuum gap. This calculation provided the first theoretical suggestion that the lattice potential of the substrate may extend significantly beyond the Å typical equilibrium spacing between atoms. It should be noted, however, that within the DFT surface science community, the emergence of charge within the vacuum region is considered a sign that the vacuum region is not large enough to create an “isolated” surface. While the calculations in question are trying to demonstrate that interaction between a substrate and film can take place across larger vacuum distances, this ignores the periodic nature of the structure. I.e., the DFT calculation represents an infinite GaAs material with repeating vacuum regions in the middle. The GaAs in the simulation is then playing the role of both the substrate and the film. It is not clear from these calculations with a single GaAs representing both the film and substrate that charge density interaction through the vacuum would persist if the film and substrate were truly isolated. A DFT slab calculation of this system would include separate GaAs atoms for the substrate and the film, both of which are separated from each other by substantial vacuum and, ideally, a dipole correction [43] such that the top of the film would not interact with the bottom of the substrate.
@@
A second challenge is that the raw potential maps contain more than one periodicity: the graphene lattice, the substrate lattice, and a longer-range graphene-induced reconstruction, such that
where is the free carrier screening coefficient through graphene [44]. To distinguish of the remote substrate, Ref. [33] used the following subtraction and cropping procedure: First, they subtract the short period potential for isolated graphene (), resulting in the maps of Fig. 1(c,d). Next they crop the longer range fluctuation of the relaxed slab (): for the profile above graphene/GaN (0001) (Fig. 1(d,f)), they focus on the line cut from atom 3 to atom 1, crop away the rest of the slab, then subtract a linear background. The energy difference along the path from atom 3 to atom 1 is then used as for the substrate, and this selected profile is periodically repeated in the line cut shown in Fig. 1(h).
@@
A recent analytical model by Kawasaki and Campbell further suggests that once free carrier screening from graphene is included, the remote substrate potential and the graphene potential have similar magnitude [44] (Fig. 3), and thus interference with graphene cannot be ignored. This model starts from a Morse pair interaction potential (rather than electrostatics), which includes covalent and van der Waals interactions between film and substrate, plus a van der Waals spacer layer modeled by a Lennard-Jones potential. The inclusion of free carrier screening for typical graphene carrier densities significantly attenuates the remote potential, such that for a typical graphene densities of order cm-2, the screened potential above graphene/GaAs is of order 10 meV (black curves), similar to the Lennard-Jones potential of graphene itself (dotted line). In the limit where the carrier density goes to zero, which is analogous to replacing graphene with insulating hBN (blue curve), the screened potential has magnitude meV, still much smaller than the 2.1 eV covalent bond strength for direct Ga-As bonding. The advantages of this approach are that the Morse potential describes a pairwise bonding interaction, rather than the purely electrostatic interaction computed by DFT, and is defined by simple, interpretable parameters that are well benchmarked by experiments and theory.
@@ III To what extent does graphene interact with a substrate or with a film?
Buffer graphene on Si-face SiC (0001), which forms a reconstruction, is a canonical example. A scanning tunneling microscopy (STM) image of this surface is shown in Fig. 4(e, h). Here the buffer carbon layer is not fully graphitic. Instead, core level spectroscopy reveals that it has mixed hybridization with and character [39], strong bonding to the SiC substrate, out-of-plane dangling bonds, Å buckling in the graphene, and both in-plane and out-of-plane distortions in the first layer of SiC [39, 51, 52, 53]. This deviation from graphitic character also opens a bandgap in buffer graphene [40, 54]. The bandgap is important because it reduces the free carrier screening of the substrate potential, compared to the screening from metallic graphene. Thus, the remote potential through a semiconducting buffer graphene is expected to be larger than the remote substrate potential through metallic graphene. However, the buckled buffer layer, with its dangling bonds, may introduce larger local potential fluctuations than flat graphene, potentially dominating over any remote field effects from the substrate.
@@
The large structural distortion of buffer graphene has a significant impact on adatom sticking and film growth. Figure 5(c) shows an STM image of Co atoms deposited on a surface that contains regions of buffer graphene and regions with a second “epitaxial” layer of graphene on top of the buffer [46]. Co clusters stick with higher probability to buffer graphene, than to epitaxial graphene. This preferred sticking on buffer graphene likely arises from the dangling bonds in hybridized buffer graphene, compared to the more idealized character of epitaxial graphene. Presumably this change in bonding character would be relevant for the growth of remote epitaxial films on SiC buffer layer, indicating a large role for reconstruction in these systems.
@@
Graphene on Ge (110) also reconstructs, into an apparent “” pattern (Fig. 4(f, i)). This reconstruction is characterized by lateral and out of plane reconstructions of the top two Ge layers, with minimal distortions in the graphene itself [47, 55, 56]. Here the graphene retains its semimetallic character. Note while this reconstruction is commonly referred to as “”, its true periodicity is defined by , where are the vectors of the surface reconstruction and are the Ge(110) surface unit cell vectors [47].
@@
Finally, even graphene on metallic surfaces induces reconstructions. Fig. 4(g, j) shows a STM image of graphene on Ir (111). The 10% lattice mismatch between graphene and Ir produces a supercell with nm periodicity, corresponding to Ir unit cells and graphene unit cells. Photoemission and scanning probe measurements reveal that the graphene is buckled by Å, due to the periodic modulation in Ir-C atomic alignment, which changes the Ir - C hybridization strength [57, 58, 48]. Moiré-driven reconstructions have also been observed on gr/Ru(0001)[59, 60, 61], gr/Rh(111)[62, 63], gr/Ni(111)[64, 65], gr/Pt(111)[66], and gr/Cu(111)[67, 68, 69] with interaction strengths spanning from strong chemisorption (e.g., Ni, Rh) to weak van der Waals bonding (e.g., Cu, Pt) [58, 57].
@@
Epitaxial ordering on graphene/metal moiré templates lies in apparent contrast to the polar bonding hypothesis for remote epitaxy. Recall that the electrostatic potential analysis by Kong et. al. suggested the remote potential fluctuations should be strongest for materials with polar bonding [33]. In contrast, Ir (111) has metallic bonding. For Ir on graphene/Ir (111) the ordering is thought to be driven by the reactivity of Ir to the HCP sites within the moiré cell [48], rather than remote epitaxy through transparent graphene.
The experiments on moire epitaxy raise the question of which lattice would the film be interacting with: the bulk substrate or the graphene induced reconstruction (or both)? While the above experiments clearly demonstrate that long-wavelength (few nanometer) moire potentials can template ordered arrays of adatom clusters, it is less clear how that long wavelength periodicity might control crystallographic orientations of films or clusters, since the characteristic spacing between atoms is few Angstrom. From the perspective of nearest-neighbor ordering in a growing crystal, several nanometer wavelength moire potentials may be too long to dictate crystallographic ordering. On the other hand, for intermediate size reconstructions, where the reconstruction wavelength is just a few multiples of the native crystal lattice constant, it is plausible that these reconstructions could be important to determining crystallographic orientation, especially if epitaxy occurs via a moderate wavelength coincident site lattice relationship rather than a simple 1:1 film:substrate relationship. Note that local variations in graphene/substrate stacking order within a moire cell can control the orientation of small molecules [60], suggesting that when the characteristic length scale for orientational ordering ( nm for molecules) matches the wavelength of the potential fluctuation, then reconstructions or moires can indeed induce that type of ordering.
@@ IV How do graphene/substrate interactions tune adatom kinetics?
Just as the underlying substrate is expected to tune the remote lattice potential, the substrate is also expected to tune adatom diffusion on graphene-covered surfaces. Growth kinetics have been underexplored in the context of remote epitaxy, but have been considered more extensively in the surface science literature.Both the presence of graphene and the identity of the underlying substrate have significant impacts on the adsorption energies, sticking coefficients, and surface diffusion of adatoms on graphene-covered surfaces. Regarding adsorption and sticking, Manzo et al.,[73] showed that the sticking of Ga and As on graphene/Ge (111) is significantly reduced compared to sticking on bare Ge(111). Poon et. al.[46] showed that Co adatom clusters stick with higher probability to buffer graphene than to epitaxial graphene on SiC, where buffer graphene is more reactive due to is more -like character compared to -like character of epitaxial graphene (Fig. 5(c)).
Fig. 6 summarizes the DFT computed surface diffusion barrier heights for several metal adatoms on graphene-covered metal surfaces [74, 75, 76, 77, 78, 79, 80, 81]. In Fig. 6(a) we plot versus the measured or computed topographical corrugation height of graphene on these surfaces, since this corrugation is a proxy for the graphene-substrate interaction strength [58]. As the substrate is changed, for Au adatoms systematically increases going from bare graphene ( Å)[82] to graphene/Cu (111) ( Å),[83] to buffer graphene/SiC ( Å)[82] to graphene/Ru(0001) ( Å).[84] For a fixed graphene/Ru(0001) substrate and varying the adatom (Fig. 6(b)), the barrier height also changes significantly and follows a general trend of valence electron count of the adatom. These data suggest that kinetics, which are tunable via the graphene/substrate interface, may play a significant role in tuning the growth mechanisms and structure of films grown on graphene-covered surfaces, in addition to the thermodynamic conditions discussed in the previous sections. “Remote” tuned diffusion barriers go beyond the typical kinetic tuning knobs of growth rate and sample temperature.
@@
For epitaxy on graphene-covered surfaces, the kinetic regimes are not well known, although one limiting case has been studied. For transferred graphene-covered surfaces, where pinholes are the primary defect, Manzo et. al. showed that when the diffusion length is greater than the spacing between pinholes , nucleation occurs preferentially at pinholes where there is direct bonding to the substrate, rather than a true “remote” epitaxy mechanism [30]. Note that where is the diffusion time. This selective nucleation is followed by lateral coalescence of a continuous film. Since the pinholes are typically only a few nanometers in diameter, films grown by the pinhole mechanism can be exfoliated, and thus share many of the macroscopic hallmarks that are expected for films that grow by a remote epitaxy mechanism. Further studies are required to investigate how tuning surface diffusion, and tuning both the identity and concentration of different types of defects, tunes the growth mode on graphene-covered surfaces. While a pinhole mechanism dominates when there is a high concentration of pinholes or larger openings in graphene [73, 87, 17, 88, 30, 89, 90], perhaps other sites such as graphene-covered atomic step edges or local variations in the graphene/substrate registry could be the nucleation sites for a true remote epitaxy mechanism (Fig. 6(c)).
V Comparison with experiment
Our analysis of the literature suggests that the remote substrate potential is not as strong as often assumed. Rather, its typical magnitude (tens of meV) is similar to the direct potential from graphene and the potential from graphene-induced interfacial reconstructions. What explains the existing experiments on remote epitaxy?
We argue there are two competing mechanisms, pinhole-seeded epitaxy [30] and van der Waals epitaxy to graphene[12], that compete with remote epitaxy and are often difficult to distinguish from a true remote mechanism. Pinhole epitaxy often occurs because defects are strong ( eV) local perturbations to Eq 2 compared to the native meV fluctuations above clean graphene (Fig. 4). Most transferred graphene platforms are not clean enough [30, 32, 31]. VdW epitaxy to graphene occurs in systems with weak remote substrate potential, where lattice matching to graphene competes [12]. It is difficult to distinguish from “remote” epitaxy if there are serial epitaxial relationships between film and graphene, and graphene and substrate. We briefly survey the main experimental observations, explain how many of these observations can be explained by these alternate mechanisms, and conclude with one experiment that cannot be explained by pinholes or epitaxy to graphene [15].
Lack of observed pinholes by microscopy. A lack of observed pinholes in graphene, either before[33, 1] or after[91] epitaxial growth, is often invoked as evidence to rule out pinhole seeded lateral epitaxy, in favor of remote epitaxy. The challenge with the before-growth observations is that they typically overlook a critical step: annealing the graphene/substrate interface. Previous experiments[30, 32] demonstrate that the standard wet and semi-dry graphene transfer procedures leave trapped contaminants and native oxides at transferred graphene/III-V substrate interfaces. Upon annealing to temperatures required for III-V film epitaxy, these contaminants volatilize to create pinholes in the graphene with densities of m-2, three orders of magnitude over thresholds needed to observe true intrinsic remote epitaxy [30] (Fig. 7a). Most graphene characterization for remote epitaxy is performed this critical annealing step, and thus cannot rule out pinholes. Pinholes down to the substrate often serve as the reactive sites for selective area epitaxy, followed by lateral coalescence of continuous films.
The challenge with post-growth defect analysis is that buried defects in graphene are difficult to resolve. Moreover, sufficient statistics are required to rule out their impact on the growth mechanism. The lack of an observed graphene pinhole by cross sectional STEM of an individual BaTiO3 island grown on graphene/SrTiO3 (001) was cited as evidence favoring remote epitaxy [91]. However, greater statistics and larger field of view imaging are required to understand if the alignment of the BaTiO3 island was due to “remote” interactions, random polydispersity, or due to features outside the field of view of individual images. Cross sectional TEM is inherently challenging for such analyses, since one of the surface directions is viewed in projection, through a finite thicknesses (typically nm) along the electron beam direction. Sophisticated techniques like ptychography are required to resolve such depth dependent information. These advanced techniques, which inherently focus on small fields of view, make it inherently difficult to generate sufficient statistics.
Epitaxy “through” multiple layers of graphene. Epitaxial films have been reported for growth on few-layer graphene on III-V and oxide substrates [33]. For these systems it is assumed that pinholes are less likely to dominate the growth mechanisms on multilayer graphene, since such pinholes would need to align vertically. However, if pinholes are created by desorption of native oxides or contaminants at the few layer graphene / substrate interface, these pinholes would be aligned [30]. Moreover, the significant attenuation due to screening suggests that epitaxy through more than two or three graphene layers is unlikely [44].
Ability to exfoliate a membrane. The ability to exfoliate a membrane from a (graphene-covered) substrate is often invoked as evidence for remote epitaxy [1, 33, 92, 93]. However, films seeded at pinholes can also be exfoliated [30]. Moreover, metal stressor layers deposited on top of the membrane are also commonly used to aid in exfoliation: presumably such stressors would not be needed if the film / graphene / substrate interfaces have sufficiently low defect density and have primarily van der Waals interactions. Finally, the presence of a graphene interlayer is not strictly required for exfoliation: Zhang et. al. [94] report efficient spalling of epitaxial oxide membranes from oxide substrates, by selecting materials combinations with weak interfacial bonding.
Film aligned to substrate on cleaner epitaxial graphene. This argument is often invoked for III-nitride epitaxy on graphene / SiC (0001) [42, 95, 45]. Here, the graphene is grown epitaxially on SiC (0001), and thus has lower defect and contaminant density than transferred graphene platforms. The challenge is that it is difficult to distinguish “remote” epitaxy from serial van der Waals epitaxy. Graphene ( Å) grows epitaxially on SiC ( Å) in a commensurate site lattice condition where a graphene superstructure nearly matches a supercell of SiC. AlN ( Å) has nearly identical lattice parameter as SiC, and thus a similar rotated commensurate site lattice relationship is expected between AlN and graphene. Thus for an all epitaxial AlN/graphene/SiC (0001) heterostructure, it is difficult to distinguish whether the structure forms via “remote” AlN-SiC interactions through graphene, or via a series of direct epitaxial interactions between AlN / graphene, and then graphene / SiC (Fig. 7b). Both mechanisms would produce the same oriented structure.
Excluding pinholes and mutual vdW epitaxy: rotated GdPtSb growth on graphene/sapphire. We end with one observation that, to our knowledge, cannot be explained by pinholes or vdW epitaxy. Du et. al.[15] found that GdPtSb films grow on polycrystalline graphene / Al2O3 (0001) with a rotation compared to direct epitaxy of GdPtSb on sapphire (Fig. 8). This rotated ordering cannot be explained by pinholes (which produce a orientation) or vdW epitaxy to the polycrystalline graphene (which would produce random in-plane oriented GdPtSb). Importantly, the new epitaxial orientation for GdPtSb on graphene/sapphire is a long-range signature that does not rely on local microscopy to exclude pinhole or vdW mechanisms.
VI Outlook